Parts (A) through (D) are four different CVP analyses, but some information is missing from each analysis. Each part contains a separate analysis.

Fill in the unknowns for each part.

Part Sales Fixed Costs Variable Costs Total Costs Contribution Margin % Operating Income
A $3,100 $1,310 $ $ 65 % $
B 4,000 3,000 4,000 %
C 6,200 930 % 310
D 460 1,110 % 3,49

*Need to fill in the missing information*
Answer #1

Total cost=Fixed costs+Variable costs

Contribution margin=Sales-Variable cost  
Contribution margin ratio=Contribution margin/Sales

Contribution margin-Fixed costs=Operating income

Sales-Total costs=Operating income

Part Sales Fixed costs Variable costs Total costs Contribution margin% Operating income
A 3100 1310 3100*(1-0.65)=$1085 (1310+1085)=$2395 65% (3100-2395)=$705
B 4000 3000 (4000-3000)=1000 4000 (4000-1000)/4000=75% (4000-4000)=$0
C 6200 930 (5890-930)=$4960 (6200-310)=$5890 (6200-4960)/6200=20% 310
D (1110+349)=$1459 (1110-460)=$650 460 1110 (1459-460)/1459=68.47%(Approx) 349

Operating income=Contribution margin-Fixed costs